# SL-SLAM: A robust visual-inertial SLAM based deep feature extraction and matching

### v1

@@ Abstract
This paper explores how deep learning techniques can improve visual-based SLAM performance in challenging environments. By combining deep feature extraction and deep matching methods, we introduce a versatile hybrid visual SLAM system designed to enhance adaptability in challenging scenarios, such as low-light conditions, dynamic lighting, weak-texture areas, and severe jitter. Our system supports multiple modes, including monocular, stereo, monocular-inertial, and stereo-inertial configurations. We also perform analysis how to combine visual SLAM with deep learning methods to enlighten other researches. Through extensive experiments on both public datasets and self-sampled data, we demonstrate the superiority of the SL-SLAM system over traditional approaches. The experimental results show that SL-SLAM outperforms state-of-the-art SLAM algorithms in terms of localization accuracy and tracking robustness. For the benefit of community, we make public the source code at https://github.com/zzzzxxxx111/SLslam.

### v2

@@ Abstract
In this letter, we investigates the paradigm of deep learning techniques to enhance the performance of visual-based SLAM systems, particularly in challenging environments. By leveraging deep feature extraction and matching methods, we propose a versatile hybrid visual SLAM framework aimed at improving adaptability in adverse conditions, such as low-light environments, dynamic lighting conditions, areas with weak textures, and significant camera jitter. Building on excellent learning-based algorithms of recent years, we designed from scratch a novel system that uses the same features extraction and matching approaches for all SLAM tasks. Our system supports multiple modes, including monocular, stereo, monocular-inertial, and stereo-inertial configurations, offering flexibility to address diverse real-world scenarios. Through comprehensive experiments conducted on both publicly available datasets and self-collected data, we demonstrate the superior performance of our SL-SLAM system compared to the SOTA approaches. We also conduct an in-depth analysis of the integration of visual SLAM with deep learning methods quantatively to provide insights for future research endeavors in this domain. The experimental results showcase the system’s capability in achieving higher localization accuracy and robust tracking performance. To facilitate further research and foster community collaboration, we have made the source code of our system publicly available at https://github.com/zzzzxxxx111/SLslam.